OpenStudy (anonymous):

The volume of a cube is found using the formula , where is the side length. What is the volume of a cube, in cubic feet, that is of a foot long?

zepdrix (zepdrix):

Woops your specific value and formula didn't copy/paste correctly.

zepdrix (zepdrix):

The volume of a cube is found using the formula `?`, where is the side length. What is the volume of a cube, in cubic feet, that is `?` of a foot long? Volume of a cube is \(\Large\rm V=\ell^3\) That's fine, we know that formula. What was the second piece of information supposed to be?

OpenStudy (anonymous):

3/5

zepdrix (zepdrix):

So your side length is,\[\Large\rm \ell =\frac{3}{5}\]If we plug that into our formula for volume (cubing the length),\[\Large\rm V=\left(\frac{3}{5}\right)^3\]You must apply the cube to `both` the numerator and denominator.

zepdrix (zepdrix):

What do you end up with? c:

OpenStudy (anonymous):

The volume of a cube is found using the formula l^3 , where l is the side length. What is the volume of a cube, in cubic feet, that is 3/5 of a foot long?

zepdrix (zepdrix):

If it's easier for you, you can use your calculator.\[\Large\rm \frac{3}{5}=0.6\]So we can cube that instead,\[\Large\rm V=(0.6)^3\]\[\Large\rm V=~?\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what expression is equivalent to m/n^3 ?